# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ Appendix D Omitted Results and Proofs in Section 4
$\left\|{{\bm{A}}{\bm{\theta}}^{\star}}\right\|^{2}=$ $\left({{\bm{\theta}}^{\star}}\right)^{\intercal}{\bm{A}}^{\intercal}{\bm{A}}{\bm{\theta}}^{\star}$
$\leq$ $2\left({\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}+\gamma\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}}\right)$
$\leq$ $2(1+\gamma)\sup_{s\in{\mathcal{S}}}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}^{2}$
@@
By Proposition B.3 with ${\bm{\eta}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$ satisfying $\eta(\tilde{s})=\nu$ for all $\tilde{s}\in{\mathcal{S}}$, where $\nu=\left({K+1}\right)^{-1}\sum_{k=0}^{K}\delta_{x_{k}}$ is the discrete uniform distribution, we can derive that, for any $r\in[0,1]$ and $s^{\prime}\in{\mathcal{S}}$, it holds that
@@
where we used the orthogonal decomposition (Proposition B.2) and an upper bound for $\ell_{2}\left({(b_{r,\gamma})_{\#}(\nu),\nu}\right)$ (Lemma E.4).
@@
$\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)=$ ${\mathbb{E}}\left[{\left\|{{\bm{A}}{\bm{\theta}}^{\star}-{\bm{b}}}\right\|^{2}}\right]\leq 2\left({{\bm{\theta}}^{\star}}\right)^{\intercal}{\mathbb{E}}\left[{{\bm{A}}^{\intercal}{\bm{A}}}\right]{\bm{\theta}}^{\star}+2{\mathbb{E}}\left[{{\bm{b}}^{\intercal}{\bm{b}}}\right].$
@@
$\left({{\bm{\theta}}^{\star}}\right)^{\intercal}{\mathbb{E}}\left[{{\bm{A}}^{\intercal}{\bm{A}}}\right]{\bm{\theta}}^{\star}\leq 2(1+\gamma)\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}^{2}.$
And by Lemma E.4,
${\mathbb{E}}\left[{{\bm{b}}^{\intercal}{\bm{b}}}\right]\leq$ $9K(1-\gamma)^{2}.$
@@
For simplicity, we use the same abbreviations as in Appendix D.2. As in the proof of [Lemma 2 Samsonov et al., 2024b], we only need to show that, for any $p\in{\mathbb{N}}$, $\alpha\in\left({0,(1-\sqrt{\gamma})/(38p)}\right)$, it holds that
${\mathbb{E}}\left\{{\left[{\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)^{\intercal}\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)}\right]^{p}}\right\}\preccurlyeq{\bm{I}}_{dK}-\frac{1}{2}\alpha p(1-\gamma){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}\left({\preccurlyeq\left({1-\frac{1}{2}\alpha p(1-\gamma)\lambda_{\min}}\right){\bm{I}}_{dK}}\right).$
Let ${\bm{B}}:={\bm{A}}+{\bm{A}}^{\intercal}-\alpha{\bm{A}}^{\intercal}{\bm{A}}$ which satisfies $\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)^{\intercal}\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)={\bm{I}}_{dK}-\alpha{\bm{B}}$, we need to bound ${\mathbb{E}}\left[{\left({{\bm{I}}_{dK}-\alpha{\bm{B}}}\right)^{p}}\right]$, it suffices to show that
@@
To show ${\mathbb{E}}\left[{{\bm{B}}}\right]\succcurlyeq(1-\sqrt{\gamma}){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}$, we first show that ${\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\intercal}}\right]\succcurlyeq 2(1-\sqrt{\gamma}){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}$, which is equivalent to $\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\intercal}}\right]\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}\succcurlyeq 2(1-\sqrt{\gamma})$, where
${\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\intercal}}\right]$ $=2\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)-{\mathbb{E}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]-{\mathbb{E}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]^{\intercal}.$
Then, for any ${\bm{w}}\in{\mathbb{R}}^{Kd}$ with $\left\|{{\bm{w}}}\right\|=1$,
${\bm{w}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\intercal}}\right]\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\bm{w}}$
$\qquad=2-2{\bm{w}}^{\intercal}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\mathbb{E}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\bm{w}}$
$\qquad\geq 2-2\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|$
@@
Next, we give an upper bound for ${\mathbb{E}}\left[{{\bm{A}}^{\intercal}{\bm{A}}}\right]$, we need to compute the following terms: by Lemma A.4,
$\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]^{2}=$ ${\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)$ (28)
$=$ $\left\|{{\bm{\phi}}(s)}\right\|^{2}{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)$
@@
${\mathbb{E}}\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]^{2}\preccurlyeq$ ${\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$
And by Lemma A.4 and Lemma E.3,
$\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]^{\intercal}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]$
$\qquad=\left({{\bm{C}}^{-\intercal}\tilde{{\bm{G}}}^{\intercal}(r){\bm{C}}^{\intercal}{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s)^{\intercal}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)$
$\qquad=\left\|{{\bm{\phi}}(s)}\right\|^{2}\left({{\bm{C}}^{-\intercal}\tilde{{\bm{G}}}^{\intercal}(r){\bm{C}}^{\intercal}{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right)$
$\qquad\preccurlyeq\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|^{2}{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right)$
$\qquad\preccurlyeq\gamma{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right).$
To summarize, by the basic inequality $\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)^{\intercal}\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)\preccurlyeq 2\left({{\bm{B}}_{1}^{\intercal}{\bm{B}}_{1}+{\bm{B}}_{2}^{\intercal}{\bm{B}}_{2}}\right)$, we have
${\bm{A}}^{\intercal}{\bm{A}}\preccurlyeq$ $2{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}+\gamma{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right),$ (29)
@@
${\mathbb{E}}\left[{{\bm{A}}^{\intercal}{\bm{A}}}\right]\preccurlyeq$ $2\left({1+\gamma}\right){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$ (30)
Combine these together, we obtain
@@
$\left\|{{\bm{A}}^{\intercal}{\bm{A}}}\right\|\leq$ $2\left\|{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}+\gamma{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right\|$ (31)
$\leq$ $2\left\|{{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}+\gamma{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\intercal}}\right\|$
@@
$\left\|{{\bm{B}}}\right\|=$ $\left\|{{\bm{A}}+{\bm{A}}^{\intercal}-\alpha{\bm{A}}^{\intercal}{\bm{A}}}\right\|$ (32)
$\leq$ $2\left\|{{\bm{A}}}\right\|+\alpha\left\|{{\bm{A}}^{\intercal}{\bm{A}}}\right\|$
@@
${\bm{B}}^{2}=$ $\left({{\bm{A}}+{\bm{A}}^{\intercal}-\alpha{\bm{A}}^{\intercal}{\bm{A}}}\right)^{2}$
@@
where we used the fact that $\left({{\bm{B}}_{1}+{\bm{B}}_{2}}\right)^{2}\preccurlyeq(1+\beta){\bm{B}}_{1}^{2}+(1+\beta^{-1}){\bm{B}}_{2}^{2}$ for any symmetric matrices ${\bm{B}}_{1},{\bm{B}}_{2}$, since $\beta{\bm{B}}_{1}^{2}+\beta^{-1}{\bm{B}}_{2}^{2}-{\bm{B}}_{1}{\bm{B}}_{2}-{\bm{B}}_{2}{\bm{B}}_{1}=\left({\sqrt{\beta}{\bm{B}}_{1}-\sqrt{\beta^{-1}}{\bm{B}}_{2}}\right)^{2}\succcurlyeq\bm{0}$, $\beta\in(0,1)$ to be determined; and the fact that ${\bm{A}}^{2}+\left({{\bm{A}}^{\intercal}}\right)^{2}\preccurlyeq{\bm{A}}^{\intercal}{\bm{A}}+{\bm{A}}{\bm{A}}^{\intercal}$ since the square of the skew-symmetric matrix is negative semi-definite $\left({{\bm{A}}-{\bm{A}}^{\intercal}}\right)^{2}\preccurlyeq\bm{0}$. By Eqn. (31) and Eqn. (30), we have
$\left\|{{\bm{A}}^{\intercal}{\bm{A}}}\right\|\leq 2(1+\gamma),$
${\mathbb{E}}\left[{{\bm{A}}^{\intercal}{\bm{A}}}\right]\preccurlyeq$ $2\left({1+\gamma}\right){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}},$ (33)
@@
$\alpha^{2}{\mathbb{E}}\left[{\left({{\bm{A}}^{\intercal}{\bm{A}}}\right)^{2}}\right]\preccurlyeq 4\alpha^{2}(1+\gamma)^{2}{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}\preccurlyeq(1-\sqrt{\gamma})^{2}{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$ (34)
As for ${\mathbb{E}}\left[{{\bm{A}}{\bm{A}}^{\intercal}}\right]$, by the basic inequality $\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)^{\intercal}\preccurlyeq 2\left({{\bm{B}}_{1}{\bm{B}}_{1}^{\intercal}+{\bm{B}}_{2}{\bm{B}}_{2}^{\intercal}}\right)$, we have
${\bm{A}}{\bm{A}}^{\intercal}=$$\left\{{\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]-\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]}\right\}$
$\cdot\left\{{\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\intercal}}\right)}\right]-\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right)}\right]}\right\}^{\intercal}$

### v2

@@ Appendix E Omitted Results and Proofs in Section 4
$\left\|{{\bm{A}}{\bm{\theta}}^{\star}}\right\|^{2}=$ $\left({{\bm{\theta}}^{\star}}\right)^{\top}{\bm{A}}^{\top}{\bm{A}}{\bm{\theta}}^{\star}$
$\leq$ $2\left({\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)}+\gamma\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\top}}\right)}}\right)$
$\leq$ $2(1+\gamma)\sup_{s\in{\mathcal{S}}}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)}^{2}$
@@
By Proposition C.3 with ${\bm{\eta}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$ satisfying $\eta(\tilde{s})=\nu$ for all $\tilde{s}\in{\mathcal{S}}$, where $\nu=\left({K+1}\right)^{-1}\sum_{k=0}^{K}\delta_{x_{k}}$ is the discrete uniform distribution, we can derive that, for any $r\in[0,1]$ and $s^{\prime}\in{\mathcal{S}}$, it holds that
@@
where we used the orthogonal decomposition (Proposition C.2) and an upper bound for $\ell_{2}\left({(b_{r,\gamma})_{\#}(\nu),\nu}\right)$ (Lemma G.4).
@@
$\operatorname{tr}\left({{\bm{\Sigma}}_{e}}\right)=$ ${\mathbb{E}}\left[{\left\|{{\bm{A}}{\bm{\theta}}^{\star}-{\bm{b}}}\right\|^{2}}\right]\leq 2\left({{\bm{\theta}}^{\star}}\right)^{\top}{\mathbb{E}}\left[{{\bm{A}}^{\top}{\bm{A}}}\right]{\bm{\theta}}^{\star}+2{\mathbb{E}}\left[{{\bm{b}}^{\top}{\bm{b}}}\right].$
@@
$\left({{\bm{\theta}}^{\star}}\right)^{\top}{\mathbb{E}}\left[{{\bm{A}}^{\top}{\bm{A}}}\right]{\bm{\theta}}^{\star}\leq 2(1+\gamma)\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}^{2}.$
And by Lemma G.4,
${\mathbb{E}}\left[{{\bm{b}}^{\top}{\bm{b}}}\right]\leq$ $9K(1-\gamma)^{2}.$
@@
For simplicity, we use the same abbreviations as in Appendix E.2. As in the proof of [Lemma 2 54], we only need to show that, for any $p\in{\mathbb{N}}$, $\alpha\in\left({0,(1-\sqrt{\gamma})/(38p)}\right)$, it holds that
${\mathbb{E}}\left\{{\left[{\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)^{\top}\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)}\right]^{p}}\right\}\preccurlyeq{\bm{I}}_{dK}-\frac{1}{2}\alpha p(1-\gamma){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}\left({\preccurlyeq\left({1-\frac{1}{2}\alpha p(1-\gamma)\lambda_{\min}}\right){\bm{I}}_{dK}}\right).$
Let ${\bm{B}}:={\bm{A}}+{\bm{A}}^{\top}-\alpha{\bm{A}}^{\top}{\bm{A}}$ which satisfies $\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)^{\top}\left({{\bm{I}}_{dK}-\alpha{\bm{A}}}\right)={\bm{I}}_{dK}-\alpha{\bm{B}}$. To give an upper bound of ${\mathbb{E}}\left[{\left({{\bm{I}}_{dK}-\alpha{\bm{B}}}\right)^{p}}\right]$, it suffices to show that
@@
To show ${\mathbb{E}}\left[{{\bm{B}}}\right]\succcurlyeq(1-\sqrt{\gamma}){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}$, we first show that ${\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\top}}\right]\succcurlyeq 2(1-\sqrt{\gamma}){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}$, which is equivalent to $\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\top}}\right]\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}\succcurlyeq 2(1-\sqrt{\gamma})$, where
${\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\top}}\right]$ $=2\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)-{\mathbb{E}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right]-{\mathbb{E}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right]^{\top}.$
Then, for any ${\bm{\theta}}\in{\mathbb{R}}^{dK}$ with $\left\|{{\bm{\theta}}}\right\|=1$,
${\bm{\theta}}^{\top}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\mathbb{E}}\left[{{\bm{A}}+{\bm{A}}^{\top}}\right]\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\bm{\theta}}$
$\qquad=2-2{\bm{\theta}}^{\top}\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\mathbb{E}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right]\left({{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}\right)^{-\frac{1}{2}}{\bm{\theta}}$
$\qquad\geq 2-2\left\|{{\mathbb{E}}_{s,r,s^{\prime}}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Sigma}}_{{\bm{\phi}}}^{-\frac{1}{2}}}\right)}\right]}\right\|$
@@
Next, we give an upper bound for ${\mathbb{E}}\left[{{\bm{A}}^{\top}{\bm{A}}}\right]$, we need to compute the following terms: by Lemma A.4,
$\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)}\right]^{2}=$ ${\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)$ (31)
$=$ $\left\|{{\bm{\phi}}(s)}\right\|^{2}{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)$
@@
${\mathbb{E}}\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)}\right]^{2}\preccurlyeq$ ${\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$
And by Lemma A.4 and Lemma G.3,
$\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right]^{\top}\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right]$
$\qquad=\left({{\bm{C}}^{-\top}\tilde{{\bm{G}}}^{\top}(r){\bm{C}}^{\top}{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s)^{\top}{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)$
$\qquad=\left\|{{\bm{\phi}}(s)}\right\|^{2}\left({{\bm{C}}^{-\top}\tilde{{\bm{G}}}^{\top}(r){\bm{C}}^{\top}{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\top}}\right)$
$\qquad\preccurlyeq\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|^{2}{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\top}}\right)$
$\qquad\preccurlyeq\gamma{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\top}}\right).$
To summarize, by the basic inequality $\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)^{\top}\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)\preccurlyeq 2\left({{\bm{B}}_{1}^{\top}{\bm{B}}_{1}+{\bm{B}}_{2}^{\top}{\bm{B}}_{2}}\right)$, we have
${\bm{A}}^{\top}{\bm{A}}\preccurlyeq$ $2{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}+\gamma{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\top}}\right),$ (32)
@@
${\mathbb{E}}\left[{{\bm{A}}^{\top}{\bm{A}}}\right]\preccurlyeq$ $2\left({1+\gamma}\right){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$ (33)
Combining these together, we obtain
@@
$\left\|{{\bm{A}}^{\top}{\bm{A}}}\right\|\leq$ $2\left\|{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}+\gamma{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right\|$ (34)
$\leq$ $2\left\|{{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}+\gamma{\bm{\phi}}(s^{\prime}){\bm{\phi}}(s^{\prime})^{\top}}\right\|$
@@
$\left\|{{\bm{B}}}\right\|=$ $\left\|{{\bm{A}}+{\bm{A}}^{\top}-\alpha{\bm{A}}^{\top}{\bm{A}}}\right\|$ (35)
$\leq$ $2\left\|{{\bm{A}}}\right\|+\alpha\left\|{{\bm{A}}^{\top}{\bm{A}}}\right\|$
@@
${\bm{B}}^{2}=$ $\left({{\bm{A}}+{\bm{A}}^{\top}-\alpha{\bm{A}}^{\top}{\bm{A}}}\right)^{2}$
@@
where we used the fact that $\left({{\bm{B}}_{1}+{\bm{B}}_{2}}\right)^{2}\preccurlyeq(1+\beta){\bm{B}}_{1}^{2}+(1+\beta^{-1}){\bm{B}}_{2}^{2}$ for any symmetric matrices ${\bm{B}}_{1},{\bm{B}}_{2}$, since $\beta{\bm{B}}_{1}^{2}+\beta^{-1}{\bm{B}}_{2}^{2}-{\bm{B}}_{1}{\bm{B}}_{2}-{\bm{B}}_{2}{\bm{B}}_{1}=\left({\sqrt{\beta}{\bm{B}}_{1}-\sqrt{\beta^{-1}}{\bm{B}}_{2}}\right)^{2}\succcurlyeq\bm{0}$, $\beta\in(0,1)$ to be determined; and the fact that ${\bm{A}}^{2}+\left({{\bm{A}}^{\top}}\right)^{2}\preccurlyeq{\bm{A}}^{\top}{\bm{A}}+{\bm{A}}{\bm{A}}^{\top}$ since the square of the skew-symmetric matrix is negative semi-definite $\left({{\bm{A}}-{\bm{A}}^{\top}}\right)^{2}\preccurlyeq\bm{0}$. By Eqn. (34) and Eqn. (33), we have
$\left\|{{\bm{A}}^{\top}{\bm{A}}}\right\|\leq 2(1+\gamma),$
${\mathbb{E}}\left[{{\bm{A}}^{\top}{\bm{A}}}\right]\preccurlyeq$ $2\left({1+\gamma}\right){\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}},$ (36)
@@
$\alpha^{2}{\mathbb{E}}\left[{\left({{\bm{A}}^{\top}{\bm{A}}}\right)^{2}}\right]\preccurlyeq 4\alpha^{2}(1+\gamma)^{2}{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}\preccurlyeq(1-\sqrt{\gamma})^{2}{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}.$ (37)
As for ${\mathbb{E}}\left[{{\bm{A}}{\bm{A}}^{\top}}\right]$, by the basic inequality $\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)\left({{\bm{B}}_{1}-{\bm{B}}_{2}}\right)^{\top}\preccurlyeq 2\left({{\bm{B}}_{1}{\bm{B}}_{1}^{\top}+{\bm{B}}_{2}{\bm{B}}_{2}^{\top}}\right)$, we have
${\bm{A}}{\bm{A}}^{\top}=$$\left\{{\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)}\right]-\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right]}\right\}$
$\cdot\left\{{\left[{{\bm{I}}_{K}\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right)}\right]-\left[{\left({{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right)\otimes\left({{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right)}\right]}\right\}^{\top}$